Companies with excess cash often employ share repurchase plans in place of or along with cash dividends. Share repurchase plans can help investors liquidate their holdings by selling their stock to the issuing company and earning from capital gains.

Consider the case of St. Sebastian Company:

St. Sebastian Company has forecasted a net income of $5,100,000 for this year. Its common stock currently trades at $20 per share, and the company currently has 790,000 shares of common stock outstanding. It has sufficient funds available to pay a cash dividend, but many of its investors don't like the additional tax liability to which the dividend income subjects them.

As a result, St. Sebastian’s management is considering making a share repurchase transaction in which it would buy back 75,000 shares of its outstanding shares in the open market by paying the current market share price. Assume that the repurchase transaction will have no effect on either the company's net income or its price-to-earnings (P/E) ratio. What is St. Sebastian's expected stock price after the stock repurchase transaction? (Note: Round your intermediate calculation to two decimal places. Round your answer to two decimal places.)

$19.89 per share

$22.10 per share

$24.31 per share

$23.21 per share

Which of these factors are considered an advantage of a stock repurchase? Check all that apply.

When a firm distributes cash by repurchasing stock, stockholders have the option to either sell or not sell stock.

Repurchases can be used to produce large-scale changes in capital structure.

Stockholders who sell their stock back to the company might claim that they were not made fully aware of all implications of the repurchase.
Answer #1

NET INCOME = $5,100,000, PRICE PER COMMON SHARE = $20

ORIGINAL SHARES = 790,000

ORIGINAL EARNING PER SHARE (EPS) = NET INCOME / ORIGINAL SHARES = $5,100,000 / 790,000 = $6.46 / SHARE

ORIGINAL PRICE TO EARNING RATIO P/E = PRICE PER COMMON SHARE / ORIGINAL EPS = $20 / $6.46 = 3.10

AS REQUIRED, P/E = CONSTANT @3.10, NET INCOME = CONSTANT @ $5,100,000

COMMON SHARES TO BUY BACK = 75,000

THEREFORE OUTSTANDING COMMON SHARES = 790,000 - 75,000 = 715,000

NEW EPS = $5,100,000 / 715,000 = 7.13

P/E = BUYBACK PRICE / OUTSTANDING COMMON SHARES

APPLYING VALUES AS ABOVE,

3.10 = BUYBACK PRICE / 7.13

THEREFORE, BUYBACK PRICE = 3.1 X 7.13 = $22.10 (OPTION II)